OpenStudy (faman39):

A number is increased by 90%, or360 . What is the original number? 189 270 400 3,600

OpenStudy (anonymous):

Do you mean when the number is increased by 90%, it is increased by 360?

OpenStudy (anonymous):

If that's the case, then let x be the number. x (90%) = 360 Solve x.

OpenStudy (unklerhaukus):

\[90\%=360\] \[10\%=\frac{360}9\]\[100\%=10\times\frac{360}9\]

OpenStudy (faman39):

Thank u so much guys! u guys r marvelous!!!!!!!!!!! :)

OpenStudy (anonymous):

Welcome :)

OpenStudy (unklerhaukus):

anytime

OpenStudy (anonymous):

400